# On the strong DR/DZ equivalence conjecture

### v1

@@ Abstract.
We establish the Miura equivalence of two integrable systems associated to a semi-simple cohomological field theory: the double ramification hierarchy of Buryak and the Dubrovin-Zhang hierarchy. This equivalence was conjectured by Buryak and further refined by Buryak, Dubrovin, Guéré, and Rossi.
@@ 1.1. Cohomological field theories
Let $V=\langle e_{1},\dots,e_{N}\rangle$ be a finite dimensional vector space equipped with a symmetric nondegenerate bilinear form $\eta$. A cohomological field theory (CohFT) [23] is a system of linear maps
@@
(1.2) $F\coloneqq\sum\frac{\varepsilon^{2g}}{n!}\left(\int_{{\overline{\mathcal{M}}}_{g,n}}c_{g,n}(\otimes_{i=1}^{n}e_{\alpha_{i}})\prod_{i=1}^{n}\psi_{i}^{d_{i}}\right)\prod_{i=1}^{n}t^{\alpha_{i},d_{i}}$
and a partition function $Z\coloneqq\exp(\epsilon^{-2}F)$.
@@ 1.2. Dubrovin-Zhang hierarchy
One can associate to a semi-simple CohFT a Hamiltonian integrable hierarchy of hydrodynamic type. This construction was initially developed by Dubrovin and Zhang [18], and subsequently by Liu, Wang, and Zhang [24], in the framework of calibrated Frobenius manifolds. In the context of CohFTs, calibration refers to an extra homogeneity property.
@@
(1.4) $\frac{\partial^{2}F}{\partial t^{\alpha,p}\partial t^{\beta,q}}=\Omega_{\alpha,p;\beta,q}(\{w^{\gamma,d}\}),$
@@ 1.3. Buryak’s double ramification hierarchy
Buryak’s DR hierarchy [5] is the following integrable system of evolutionary PDEs:
@@ 1.4. Normal Miura equivalence
In general the Miura group acts on (Hamiltonian) integrable systems by the diffeomorphic changes of dependent variables, see e.g. [18]. More refined are the so-called normal Miura transformations that transform a tau-symmetric Hamiltonian hierarchy written in normal coordinates to a tau-symmetric Hamiltonian hierarchy written in normal coordinates.
Let $\tilde{u}^{\alpha}$ be the normal coordinates for Buryak’s double ramification hierarchy. The normal coordinates for the Dubrovin-Zhang hierarchy are $w^{\alpha}$. The following conjecture is proposed in [6, Conjecture 7.3]:
@@
For any semi-simple cohomological field theory there exists a differential polynomial $\mathcal{P}$ in coordinates $w^{\alpha}$ (and it has a very explicit description that we omit here) such that the normal Miura transformation defined as
(1.9) $w^{\alpha}\mapsto\tilde{u}^{\alpha}(w)=w^{\alpha}+\eta^{\alpha\gamma}{\partial}_{x}\{\mathcal{P},\bar{h}^{\mathrm{DZ}}_{\mu,0}\}_{K^{DZ}}$
@@ 1.5. Generalizations and approach to the proof
There is a number of generalizations of the strong DR/DZ equivalence conjecture for semi-simple CohFTs above. In particular, the condition of semi-simplicity can be dropped (but then one has to prove that the Dubrovin-Zhang integrable system is given by differential polynomials in the dependent variables), and similar constructions can be applied to the so-called partial CohFTs and F-CohFTs (in the latter case we get just a system of conservation laws rather than a Hamiltonian system). These constructions are available in [1, 14, 16], and the corresponding conjectures there state that
@@
We refer to [16, Section 4] for the precise statement and further details. In [16], following and generalizing earlier results in [6, 8], these conjectures are reduced to a system of conjectural tautological relations in the moduli space of curves known as a generalization of the $A=B$ relation. The initial $A=B$ relation itself was proposed in [8] and designed to imply Conjecture 1.1. It is convenient to let the term “$A=B$ relation” refer to the whole system of (generalized) $A=B$ relations.
In this paper we prove the $A=B$ relation as well as all its generalization from [16] in the Gorenstein quotient. In other words we prove the relations after intersection with tautological classes of the complementary dimension (Theorem 2.3). As a consequence, our result implies the entire system of conjectures for not-necessary semi-simple CohFTs, as well as partial CohFTs and F-CohFTs under the assumption that the classes involved in these structures are tautological. Since this is always the case for semi-simple CohFTs, as a corollary Theorem 2.3 implies Conjecture 1.1 (Theorem 1.2). We remark that Theorem 2.3 is the main result of this paper, and we moreover refer further to [16, Section 4] for a list of corollaries descending from it.
@@ 1.7. Acknowledgments
X. B. and S. S. were supported by the Netherlands Organization for Scientific Research. D. L. is supported by the University of Trieste, by the INFN under the national project MMNLP (APINE) section of Trieste, by the INdAM group GNSAGA, and by the PRIN project 2022 “Geometry of algebraic structures: moduli, invariants, deformations”. The authors thank A. Buryak and P. Rossi for useful discussions and collaboration on closely related topics.
@@ 2.1. Basic notation for trees
  • We say that a vertex or a (half-)edge $x$ is a descendant of a vertex or a (half-)edge $y$ if $y$ is on the unique path connecting $x$ to $v_{r}$.
@@ 2.2. Leveled stable rooted trees
    |$|\{v\in V(T)\,|\,\ell(v)\leq i\}|-1+\sum_{\begin{subarray}{c}v\in V(T)\\ \ell(v)\leq i\end{subarray}}p(v)\leq\sum_{\begin{subarray}{c}v\in V(T)\\ \ell(v)\leq i\end{subarray}}2g(v)-2+m.$
@@
The set $\mathrm{LDLSRT}_{g,n,m}$ of leveled degree-labeled stable rooted trees consists of triples $(T,p,\ell)$, where $T\in\mathrm{SRT}_{g,n,m}$, $p\in\mathcal{P}(T)$, and $\ell\in\mathcal{L}(T,p)$.
@@ 2.3. Definition of the $B$ class
$\Psi(v)\coloneqq\left(\prod_{i=1}^{|H_{+}(v)|}\frac{1}{1-a(h_{i})\psi_{i}}\right)_{p(v)}\in R^{p}({\overline{\mathcal{M}}}_{g(v),|H(v)|})\otimes_{\mathbb{Q}}Q$
@@
(2.1) $B^{m}_{g,n}\coloneqq\sum_{(T,p,\ell)\in\mathrm{SRT}(g,n,m)}(-1)^{\max\ell(V(T))}\biggl{(}\prod_{e\in E(T)}a(e)\biggr{)}(b_{T})_{*}\bigotimes_{v\in V(T)}\Psi(v)$
@@ 2.4. The conjectural $A=B$ relation revisited
Consider the moduli space
@@
of rubber stable maps to $(\mathbb{P}^{1},0,\infty)$. Let $s\colon{\overline{\mathcal{M}}}_{g}^{\sim}(\mathbb{P}^{1},a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})\to{\overline{\mathcal{M}}}_{g,n+1}$ be the projection to the source curve, and $\lambda_{g}$ the lift of the lambda class with respect to this projection. Let $t\colon{\overline{\mathcal{M}}}_{g}^{\sim}(\mathbb{P}^{1},a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})\to LM_{2g-1+n}$ be the projection to the target curve, where $LM_{2g-1+n}$ denotes the Losev-Manin space with $2g-1+n$ marked points. Let $\tilde{\psi}_{0}$ be the pull-back by $t$ of the $\psi$-class at the point 00 in Losev-Manin space. Define
@@
Note that $(A^{1}_{g,n})_{<2g}=0$ and $(A^{1}_{g,n})_{2g}=\lambda_{g}\mathrm{DR}_{g}(a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})$. Note also that for any $d\geq 2g$ the class $(A^{1}_{g,n})_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$.
Conjecture 2.1.
@@
In all three statement $\deg$ refers either to cohomological degree or to the homogenenous degree in $a_{1},\dots,a_{n}$. We use the convention that a polynomial of negative degree is the zero polynomial. Note that the third statement follows from the second statement. We have the following lemma:
@@
Conjecture 2.1 is equivalent to [16, Conjectures 1, 2, and 3].
@@
We have to show that our definition of the class $A^{1}_{g,n}$ coincides with the definition given in [16] (see also [3, Section 2.4]). To this end, we perform the intersection of the class $\tilde{\psi}_{0}$ with the moduli space of rubber maps using the technics of [17], this yields exactly the same trees in $\mathrm{SRT}_{g,n,1}$ decorated by double ramification cycles as in the original definition of $A^{1}_{g,n}$. The only difference is the combinatorial coefficient that we have to compare.
@@
For the computation of (2.3) we consider a set rigorous level structures $\mathcal{R}(T)$, that is, the set of one-to-one maps $f\colon V(T)\to\{1,\dots,|V(T)|\}$ such that $f(v_{1})>f(v_{2})$ if $v_{1}\in DV(v_{2})$. In particular, $f(v_{r})=1$. Then the coefficient implied by (2.3) is equal to

### v2

@@ Abstract.
We establish the Miura equivalence of two integrable hierarchies associated to a semi-simple cohomological field theory: the double ramification hierarchy of Buryak and the Dubrovin-Zhang hierarchy. This equivalence was conjectured by Buryak and further refined by Buryak, Dubrovin, Guéré, and Rossi.
@@ 1.1. Cohomological field theories
Let $V=\langle e_{1},\dots,e_{N}\rangle$ be a finite dimensional vector space equipped with a symmetric nondegenerate bilinear form $\eta$. The first basis vector $e_{1}$ is distinguished. A cohomological field theory (CohFT) [24] is a system of linear maps
@@
(1.2) $F\coloneqq\sum_{\begin{subarray}{c}g,n\geq 0\\ 2g-2+n>0\\ 1\leq\alpha_{1},\dots,\alpha_{n}\leq N\\ d_{1},\dots,d_{n}\geq 0\end{subarray}}\frac{\varepsilon^{2g}}{n!}\left(\int_{{\overline{\mathcal{M}}}_{g,n}}c_{g,n}(\otimes_{i=1}^{n}e_{\alpha_{i}})\prod_{i=1}^{n}\psi_{i}^{d_{i}}\right)\prod_{i=1}^{n}t^{\alpha_{i},d_{i}}$
and a partition function $Z\coloneqq\exp(\epsilon^{-2}F)$. Here $\psi_{i}$ are the tautological classes defined as first Chern classes of the cotangent bundle on the $i$-th marked point, $t^{\alpha_{i},d_{i}}$ are formal variables with $\alpha_{i}=1,\dots,N$ and $d_{i}\geq 0$.
@@ 1.2. Dubrovin-Zhang hierarchy
One can associate to a semi-simple CohFT a Hamiltonian integrable hierarchy of hydrodynamic type. This construction was initially developed by Dubrovin and Zhang [19], and subsequently by Liu, Wang, and Zhang [25], in the framework of calibrated semi-simple Frobenius manifolds. In the context of CohFTs, Frobenius manifold refers to an extra homogeneity property.
@@
It is a formal power series in $t^{\gamma,p}+\delta^{\gamma}_{1}\delta^{p}_{0}x$, $\gamma=1,\dots,N$, $p\geq 0$. The shift of $t^{1,0}$ by $x$ is needed in what follows to have a spatial variable in the construction of an integrable system. The second upper index of $w^{\alpha,d}$ can be omitted in case it is zero.
@@
(1.4) $\frac{\partial^{2}F}{\partial t^{\alpha,p}\partial t^{\beta,q}}\big{|}_{t^{1,0}\to t^{1,0}+x}=\Omega_{\alpha,p;\beta,q}(\{w^{\gamma,d}\}),$
@@ 1.3. Buryak’s double ramification hierarchy
Here $\lambda_{g}$ and $\mathrm{DR}_{g}$ are the top Chern class of the Hodge bundle and the double ramification cycle, respectively. Buryak’s DR hierarchy [6] is the following integrable system of evolutionary PDEs:
@@ 1.4. Normal Miura equivalence
In general the Miura group acts on (Hamiltonian) integrable systems by the diffeomorphic changes of dependent variables, see e.g. [19]. More refined are the so-called normal Miura transformations that transform a tau-symmetric Hamiltonian hierarchy written in normal coordinates to a tau-symmetric Hamiltonian hierarchy written in normal coordinates, see e.g. [7, Section 3].
Let $\tilde{u}^{\alpha}$ be the normal coordinates for Buryak’s double ramification hierarchy [7, Section 7]. The normal coordinates for the Dubrovin-Zhang hierarchy are $w^{\alpha}$ [19, 11]. The following conjecture is proposed in [7, Conjecture 7.3]:
@@
For any semi-simple cohomological field theory there exists a differential polynomial $\mathcal{P}$ in coordinates $w^{\alpha}$ (whose explicit expression is given in [17, Theorem 4.4], but omitted here) such that the normal Miura transformation defined as
(1.9) $w^{\alpha}\mapsto\tilde{u}^{\alpha}(w)=w^{\alpha}+\eta^{\alpha\mu}{\partial}_{x}\{\mathcal{P},\bar{h}^{\mathrm{DZ}}_{\mu,0}\}_{K^{DZ}}$
@@ 1.5. Generalizations and approach to the proof
There is a number of generalizations of the strong DR/DZ equivalence conjecture for semi-simple CohFTs above. In particular, the condition of semi-simplicity can be dropped; however, one then has to prove that the Dubrovin-Zhang integrable system is given by differential polynomials in the dependent variables. Similar constructions also apply to the so-called partial CohFTs and F-CohFTs. In the case of F-CohFTs, we get just a system of conservation laws rather than a Hamiltonian system. These constructions are available in [1, 15, 17], and the corresponding conjectures there state that
@@
We refer to [17, Section 4] for the precise statement and further details. In [17], following and generalizing earlier results in [7, 9], these conjectures are reduced to a system of conjectural tautological relations in the moduli space of curves known as a generalization of the $A=B$ relation. The initial $A=B$ relation itself was proposed in [9] and designed to imply Conjecture 1.1. It is convenient to let the term “$A=B$ relation” refer to the whole system of (generalized) $A=B$ relations.
In this paper we prove the $A=B$ relation as well as all its generalization from [17] in the Gorenstein quotient. In other words we prove the relations after intersection with tautological classes of the complementary dimension (Theorem 2.8). As a consequence, our result implies the entire system of conjectures for not-necessary semi-simple CohFTs, as well as partial CohFTs and F-CohFTs under the assumption that the classes involved in these structures are tautological. Since this is always the case for semi-simple CohFTs, as a corollary Theorem 2.8 implies Conjecture 1.1 (Theorem 1.2). We remark that Theorem 2.8 is the main result of this paper, and we moreover refer further to [17, Section 4] for a list of corollaries descending from it.
@@ 1.7. Acknowledgments
X. B. and S. S. were supported by the Netherlands Organization for Scientific Research. D. L. is supported by the University of Trieste, by the INFN under the national project MMNLP (APINE) section of Trieste, by the INdAM group GNSAGA, and by the PRIN project 2022 “Geometry of algebraic structures: moduli, invariants, deformations”. The authors thank A. Buryak, P. Rossi, and A. Sauvaget for useful discussions and collaboration on closely related topics. The authors also thank an anonymous referee for many useful remarks and questions that allowed to substantially improve the presentation.
1.8. A further development
About a month after this paper was completed and submitted, the conjecture on the master relation that we pose here, Conjecture 3.4, was proved in [5] using the virtual localization technique on the space of stable relative maps to the projective line.
This is an exciting development, and, first of all, in the framework of the present paper it provides an alternative to the argument that we give in Section 4.4. Second, by Theorem 4.5, this establishes the generalized $A=B$ relations (Conjecture 2.6), thereby establishing the most general version of the DR/DZ equivalence. As a consequence, the semi-simplicity assumption in Theorem 1.2 can be dropped. More generally, for any partial CohFT or F-CohFT, we get the following:
  • The Dubrovin–Zhang hierarchy exists; that is, the associated equations are differential polynomial, cf. [17, Theorem 4.7].
  • The Dubrovin–Zhang hierarchy is equivalent to the double ramification hierarchy, cf. [17, Theorem 4.10].
However, we strongly believe that for the integrable systems community our original proof of formally weaker Theorem 3.5 that is contained in Section 4.4 is probably more appealing and might be more useful. First of all, it does not involve any additional techniques or concepts from algebraic geometry that were not already present in the construction of the DR hierarchy. Second it is much more simple and still sufficient for the proof of Conjecture 1.1 in its original formulation. And finally, it is visibly better aligned with the methods and ideas used in the construction of the Durbovin-Zhang hierarchies in [11] in the semi-simple case and thus might be used as a tool to treat the existence of the DR form of for the DZ hierarchies as a system of their universal properties.
@@ 2.1. Basic notation for trees
  • We say that a vertex or a (half-)edge $x$ is a descendant of a vertex or a (half-)edge $y$ if $y$ is on the unique path connecting $x$ to $v_{r}$. For instance, for an edge $e$ formed by two half-edges $h_{+}\in H_{+}(T)$ and $h_{-}\in H_{-}(T)$ we assume that $e$, $h_{-}$, and $h_{+}$ are all descendants of $h_{+}$, $e$ and $h_{-}$ are both descendants of themselves and each other, and $h_{+}$ is not a descendant of either $e$ or $h_{-}$.
@@
2.2. Trees and strata
In general, stable graphs are used to represent natural strata in the moduli spaces of curves, where the vertices correspond to the irreducible components, legs to the marked points, and edges to the nodes. In our setting, to each stable rooted tree $T\in\mathrm{SRT}_{g,n,m}$ we associate the moduli space
¯$\overline{\mathcal{M}}_{T}=\prod_{v\in V(T)}\overline{\mathcal{M}}_{g\left(v\right),\left|H\left(v\right)\right|}.$
There is a canonical map, called the boundary map,
$b_{T}:\overline{\mathcal{M}}_{T}\rightarrow\overline{\mathcal{M}}_{g,n+m}$
whose image is the closure of the boundary stratum associated to the graph $T$. More details can be found in [27, Sections 0.2 and 0.3].
The stable rooted trees in $\mathrm{SRT}_{g,n,m}$ are used below to represent strata in ¯${\overline{\mathcal{M}}}_{g,n+m}$, and the extra combinatorial structure that we introduce here is used to specify the classes that we study.
@@ 2.3. Leveled stable rooted trees
Remark 2.1.
The nonnegative integer $p\left(v\right)$ represents the total degree of the cohomology classes on ¯$\overline{\mathcal{M}}_{g\left(v\right),\left|H\left(v\right)\right|}$.
@@
    (2.1) |$|\{v\in V(T)\,|\,\ell(v)\leq i\}|-1+\sum_{\begin{subarray}{c}v\in V(T)\\ \ell(v)\leq i\end{subarray}}p(v)\leq\sum_{\begin{subarray}{c}v\in V(T)\\ \ell(v)\leq i\end{subarray}}2g(v)-2+m.$
@@
The set $\mathrm{LDLSRT}_{g,n,m}$ of leveled degree-labeled stable rooted trees consists of triples $(T,p,\ell)$, where $T\in\mathrm{SRT}_{g,n,m}$, $p\in\mathcal{P}(T)$, and $\ell\in\mathcal{L}(T,p)$, and it is a finite set.
Example 2.2.
Let $T\in\mathrm{SRT}_{g,n=5,m=2}$ be the tree shown in the following picture, where $g_{0}+g_{1}+g_{2}+g_{3}=g$. Each vertex is depicted as a circle, with its genus indicated inside. The root is the vertex of genus $g_{0}$. The two frozen legs are represented by wavy lines, while the five regular legs are represented by regular lines. The half-edges of $H_{+}\left(T\right)$ are decorated with $\psi$-classes. We introduce
$p(v_{0}):=d_{1}+d_{6}+d_{7},\quad p(v_{1}):=d_{2}+d_{8},\quad p(v_{2}):=d_{3},\quad p(v_{4})=d_{4}+d_{5}.$
We suppose that $p\left(v_{i}\right)\leq 3g\left(v_{i}\right)-3+\left|H\left(v_{i}\right)\right|$, for $i=1,\dots,5$, so that $\left(T,p\right)$ is a degree-labeled stable rooted tree. A choice of a level function is represented with the dashed lines: the root is assigned level 00, the vertex $1$ lies at level $1$, and the vertices $2$ and $3$ are at level $2$. We assume that
$p\left(v_{0}\right)\leq 2g_{0},\quad{\rm and}\quad p\left(v_{0}\right)+p\left(v_{1}\right)+1\leq 2\left(g_{0}+g_{1}\right),$
so that this level function is admissible for the pair $\left(T,p\right)$. Observe that cutting the tree along the dashed line at level $i$ yields a subtree to the left, which is itself a rooted tree decorated with $\psi$-classes, and the degree of the associated cohomology class equals the left-hand side of inequality (2.1).
[…]
@@ 2.4. Definition of the $B$ class
$\Psi(v)\coloneqq\left(\prod_{i=1}^{|H_{+}(v)|}\frac{1}{1-a(h_{i})\psi_{i}}\right)_{p(v)}\in R^{p(v)}({\overline{\mathcal{M}}}_{g(v),|H(v)|})\otimes_{\mathbb{Q}}Q$
@@
(2.2) $B^{m}_{g,n}\coloneqq\sum_{(T,p,\ell)\in\mathrm{LDLSRT}(g,n,m)}(-1)^{\max\ell(V(T))}\biggl{(}\prod_{e\in E(T)}a(e)\biggr{)}(b_{T})_{*}\bigotimes_{v\in V(T)}\Psi(v)$
@@
Example 2.3.
Let $\left(T,p,\ell\right)\in\mathrm{LDLSRT}_{g,n,m}$ be as in Example 2.2. Suppose that the regular leg $\sigma_{i}$ is decorated with $\psi^{d_{i}}$, and carries the weight $a\left(\sigma_{i}\right)=a_{i}$, for $i=1,\dots,5.$ Then the coefficient of $(b_{T})_{*}\bigotimes_{v\in V(T)}\Psi(v)$ in Eq. (2.2) is given by
$\left(-1\right)^{2}\left(a_{2}+a_{3}\right)\left(a_{3}\right)\left(a_{4}+a_{5}\right).$
Remark 2.4.
A brief history of the $B$-class: the class $B^{0}$ was first introduced in [9]. In [17], this definition was generalized to $B^{m}$ for all $m\geq 0$, and reshaped using level structures. Finally, [3, Section 2.5] presents the form of the $B$-class used here. This last formulation is equivalent to the one of [17] by [3, Theorem 2.15]. Notice that in [17] the $B$-class is indexed with multi-indices $(d_{1},\dots,d_{n})$, which corresponds to extracting the coefficient of $a_{1}^{d_{1}}\cdots a_{n}^{d_{n}}$ in the above definition of $B_{g,n}^{m}$.
@@ 2.5. The conjectural $A=B$ relation revisited
Let $a_{1},\dots,a_{n}$ be a list of positive integers. Consider the moduli space
@@
of rubber stable maps to $(\mathbb{P}^{1},0,\infty)$. Let $s\colon{\overline{\mathcal{M}}}_{g}^{\sim}(\mathbb{P}^{1},a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})\to{\overline{\mathcal{M}}}_{g,n+1}$ be the projection to the source curve, and $\lambda_{g}$ the lift of the lambda class with respect to this projection. Let $t\colon{\overline{\mathcal{M}}}_{g}^{\sim}(\mathbb{P}^{1},a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})\to LM_{2g-1+n}$ be the projection to the target curve, where $LM_{2g-1+n}$ denotes the Losev-Manin space (see e.g. [3] for more details) with $2g-1+n$ marked points. Let $\tilde{\psi}_{0}$ be the pull-back by $t$ of the $\psi$-class at the point 00 in Losev-Manin space. Define
@@
Note that $(A^{1}_{g,n})_{<2g}=0$ and $(A^{1}_{g,n})_{2g}=\lambda_{g}\mathrm{DR}_{g}(a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})$. Note also that for any $d\geq 2g$ the class $(A^{1}_{g,n})_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$. This follows, for example, from the computation of the class $\tilde{\psi}_{0}$ in the proof Lemma 2.7.
Remark 2.5.
A brief history of the $A$-class: the class $A^{0}_{g,d_{1},\dots,d_{n}}\in H^{*}({\overline{\mathcal{M}}}_{g,n})$ was introduced in [8, 9] as a sum over stable trees decorated with $\lambda_{g}$-classes and DR-cycles. In a similar spirit, the class $A^{1}_{g,d_{1},\dots,d_{n}}\in H^{*}({\overline{\mathcal{M}}}_{g,n+1})$ was introduced in [17], where it was observed that $A^{0}_{g,d_{1},\dots,d_{n}}=\frac{1}{\sum a_{i}}\pi_{*}\left(A^{1}_{g,d_{1},\dots,d_{n}}\right)$, where $\pi\colon{\overline{\mathcal{M}}}_{g,n+1}\to{\overline{\mathcal{M}}}_{g,n}$ is the map that forgets the last marked point. The class $A^{1}_{g,n}$ defined in Eq. (2.4) provides a new formulation of the class $A^{1}$; as shown in Lemma 2.7 below this formulation satisfies $A^{1}_{g,d_{1},\dots,d_{n}}=\mathrm{Coeff}_{a_{1}^{d_{1}}\cdots a_{n}^{d_{n}}}A_{g,n}^{1}$, where $\mathrm{Coeff}_{a_{1}^{d_{1}}\cdots a_{n}^{d_{n}}}$ denotes the operation of extracting the coefficient of the monomial $a_{1}^{d_{1}}\cdots a_{n}^{d_{n}}$.
Conjecture 2.6 (generalized $A=B$ relations).
@@
In all three statement $\deg$ refers either to cohomological degree or to the homogenenous degree in $a_{1},\dots,a_{n}$. We use the convention that a class $\alpha\in H^{2i}(X)$ has cohomological degree $i$ instead of $2i$ and that a polynomial of negative degree is the zero polynomial. Note that the third statement follows from the second statement. We have the following lemma:
@@
Conjecture 2.6 is equivalent to [17, Conjectures 1, 2, and 3].
@@
We have to show that our definition (2.4) of the class $A^{1}_{g,n}$ coincides with the definition given in [17] (see also [3, Section 2.4] for notations matching ours). To this end, we will first perform the intersection of the class $\tilde{\psi}_{0}$ with the moduli space of rubber maps using the technics of [18], as we will show this yields exactly the same trees in $\mathrm{SRT}_{g,n,1}$ decorated by double ramification cycles as in the original definition of $A^{1}_{g,n}$. The only difference is the combinatorial coefficient that we will compare in a second time.
We express powers of $\psi_{0}$ in the Losev-Manin space $LM_{2g-1+n}$ using its expression as a boundary divisor [18, Eq. (2)] to obtain
$\psi_{0}^{d}=\sum_{\begin{subarray}{c}\chi_{1}+\cdots+\chi_{d+1}=2g-1+n\\ \chi_{1},\dots,\chi_{d+1}>0\end{subarray}}D_{\chi_{1},\dots,\chi_{d+1}}\frac{\chi_{1}}{\chi_{1}+\cdots+\chi_{d+1}}\frac{\chi_{2}}{\chi_{2}+\cdots+\chi_{d+1}}\cdots\frac{\chi_{d+1}}{\chi_{d+1}},$
where $D_{\chi_{1},\dots,\chi_{d+1}}$ denotes the boundary strata in $LM_{2g-1+m}$ composed of $d+1$ components, where we number the components such that the first component contains $\infty$, the second is attached to the first and so on, and such that the $i$th component contains $\chi_{i}$ marked points (excluding the points 00 and $\infty$ on each component). We now compute $\lambda_{g}s_{*}t^{*}\left(D_{\chi_{1},\dots,\chi_{d+1}}\right)$ using the following two observations. First, since the expression is ultimately multiplied by $\lambda_{g}$, we can consider that the prestable curves involved in $t^{*}\left(D_{\chi_{1},\dots,\chi_{d+1}}\right)$ have no loop in their dual graph. Second, by [18, Lemma 2.3], the ramifications points of total order $\chi_{i}$ lying above the $i$th component of $D_{\chi_{1},\dots,\chi_{d+1}}$ (excluding those above the nodes) must belong to a unique component, although additional unstable components may still be present. Note that by the Riemann-Huwtiz formula, the genus $g_{i}$ and the number of nodes $n_{i}$ of this stable component satisfy $\chi_{i}=2g_{i}-2+n_{i}$. As a result $\lambda_{g}s_{*}t^{*}\left(D_{\chi_{1},\dots,\chi_{d+1}}\right)$ gives a sum over trees $T\in\mathrm{SRT}_{g,n,1}$, where the contribution of each tree coincides precisely with the term $A\left(T\right)$ in the expression of the $A^{1}_{g,n}$-class from [3, Section 2.4], with combinatorial coefficient equal to $1$.
@@
For the computation of (2.4) we have to sum over the number of times the same tree $T$ appears in the computation of powers of $\tilde{\psi}_{0}$ It is easy to see that this corresponds to summing over the set of rigorous level structures $\mathcal{R}\left(T\right)$, that is, the set of one-to-one maps $f\colon V(T)\to\{1,\dots,|V(T)|\}$ such that $f(v_{1})>f(v_{2})$ if $v_{1}\in DV(v_{2})$. Then the coefficient implied by (2.4) is equal to